OpenStudy (anonymous):

Solve the system algebraically. Find the number of solutions for the system. 4x-y=-4 12x-3y=-24 a. No solution b. One solution c. Infinite number of solution d. None of these

OpenStudy (zepp):

You can always use the reduction method to solve it.

OpenStudy (zepp):

\((4x-y=-4)3 \\12x-3y=-12\)

OpenStudy (zepp):

Then simply substract them 12x-3y=-24 -12x-3y=-12 ------------- 0=-24--12 0=-24+12 0=-12 And since \(0 \ne -12\), this system has no solution.

OpenStudy (precal):

 notice the slope is 4